# Tree-level 11D supergravity amplitudes from the pure spinor worldline

### v1

@@ C.1. Compact form for the 4-point function
where $\cdots$ are expressions involving non-minimal variables, which are guaranteed to at the end of the computation. We then use this to manipulate the last line of (C.23),
$\eqalignno{&\ \ \ \ -3k_{12}\left\langle(\lambda\gamma_{ab}\lambda)\phi^{a}_{1}\phi^{b}_{2}(\phi_{3}^{e}C_{e,4}+\phi_{3}^{\alpha}C_{\alpha,4})\right\rangle=-3k_{12}\left\langle(\lambda\gamma_{ab}\lambda)\phi^{a}_{1}\phi^{b}_{2}\left(\phi_{3}^{c}{\tenbf C}_{c}+\phi_{3}^{\alpha}{\tenbf C}_{\alpha}\right)U_{4}^{(3)}\right\rangle\cr&=3k_{12}\left\langle\Big{[}\left(\phi_{3}^{c}{\tenbf C}_{c}-\phi_{3}^{\alpha}{\tenbf C}_{\alpha}\right),(\lambda\gamma_{ab}\lambda)\phi^{a}_{1}\phi^{b}_{2}\Big{]}U_{4}^{(3)}\right\rangle\cr&=-k_{12}\left\langle(M_{a,12}\phi_{3}^{a}-M_{\alpha,12}\phi_{3}^{\alpha})U_{4}^{(3)}\right\rangle\cr&\ \ \ -2k_{12}\left\langle(\lambda\gamma_{ab}\phi_{3})\phi^{a}_{1}\phi^{b}_{2}U_{4}^{(3)}\right\rangle-k_{12}\left\langle(\lambda\gamma_{ab}\phi_{1})\phi^{a}_{2}\phi^{b}_{3}U_{4}^{(3)}\right\rangle-k_{12}\left\langle(\lambda\gamma_{ab}\phi_{2})\phi^{a}_{1}\phi^{b}_{3}U_{4}^{(3)}\right\rangle.\cr&&({\hbox{C.}25})}$
@@
Therefore, it follows that (3.22)

### v2

@@ C.1. Compact form for the 4-point function
where $\cdots$ are expressions involving non-minimal variables, which are guaranteed to cancel at the end of the computation. We then use this to manipulate the last line of (C.23),
$\eqalignno{&\ \ \ \ -{3s_{12}\over 2}\left\langle(\lambda\gamma_{ab}\lambda)\phi^{a}_{1}\phi^{b}_{2}(\phi_{3}^{e}C_{e,4}+\phi_{3}^{\alpha}C_{\alpha,4})\right\rangle=-{3s_{12}\over 2}\left\langle(\lambda\gamma_{ab}\lambda)\phi^{a}_{1}\phi^{b}_{2}\left(\phi_{3}^{c}{\tenbf C}_{c}+\phi_{3}^{\alpha}{\tenbf C}_{\alpha}\right)U_{4}^{(3)}\right\rangle\cr&={3s_{12}\over 2}\left\langle\Big[\left(\phi_{3}^{c}{\tenbf C}_{c}-\phi_{3}^{\alpha}{\tenbf C}_{\alpha}\right),(\lambda\gamma_{ab}\lambda)\phi^{a}_{1}\phi^{b}_{2}\Big]U_{4}^{(3)}\right\rangle\cr&=-{s_{12}\over 2}\left\langle(M_{a,12}\phi_{3}^{a}-M_{\alpha,12}\phi_{3}^{\alpha})U_{4}^{(3)}\right\rangle\cr&\ \ \ -s_{12}\left\langle(\lambda\gamma_{ab}\phi_{3})\phi^{a}_{1}\phi^{b}_{2}U_{4}^{(3)}\right\rangle-{s_{12}\over 2}\left\langle(\lambda\gamma_{ab}\phi_{1})\phi^{a}_{2}\phi^{b}_{3}U_{4}^{(3)}\right\rangle-{s_{12}\over 2}\left\langle(\lambda\gamma_{ab}\phi_{2})\phi^{a}_{1}\phi^{b}_{3}U_{4}^{(3)}\right\rangle.\cr&&({\hbox{C.}25})}$
@@
Therefore, it follows that